# Improving absolute separability bounds for arbitrary dimensions

### v1

@@ Abstract
Sufficient analytical conditions for separability in composite quantum systems are very scarce and only known for low-dimensional cases. Here, we use linear maps and their inverses to derive powerful analytical conditions, providing tight bounds and extremal points of the set of absolutely separable states, i.e., states that remain separable under any global unitary transformation. Our analytical results apply to generic quantum states in arbitrary dimensions, and depend only on a single or very few eigenvalues of the considered state. Furthermore, we use convex geometry tools to improve the general characterization of the AS set given several non-comparable criteria. Finally, we present various conditions related to the twin problem of characterizing absolute PPT, that is, the set of quantum states that are positive under partial transposition and remain so under all unitary transformations.
footnotetext: *These authors contributed equally to this work.
@@ 1 Introduction
As it is well known, determining and quantifying entanglement has been proven to be, in general, an NP-Hard problem [1, 2]. After several decades of intense research, sufficient and necessary conditions to ensure or deny the entanglement content of a state exist only for the simplest cases [3, 4], or very restricted families of states. One of those conditions is linked to positive but not completely positive maps, but to date neither the use of a finite number of such maps nor the use of semidefinite programming computation [5, 6] can give a conclusive answer if a given state is entangled or not [7, 8]. Under these circumstances, one may relax the quest for iff conditions and search for sufficient conditions for separability.
@@
Here, we wish to draw attention to the use of not completely positive maps and their inverses, as sufficient separability criteria [9, 10, 11]. In particular, we focus here on the concept of absolute separability (AS), that is, those states whose separability is preserved under all unitary transformations. Characterizing the set of AS is relevant for various reasons. Asides of its importance for the general characterization of the set of quantum states, AS states become pertinent in the resource theory of entanglement, as they represent free states, while entangled states act as a resource. Since unitary transformations are the most common method to generate entanglement, e.g., through quantum circuits or quenches, if a state is certified AS it cannot lead to any entanglement under unitary transformations [12]. Moreover, AS states depend only on their spectrum (eigenvalues), and their characterization is significantly less demanding than performing full quantum state tomography [13, 14]. For a quantum state of dimension $D=N\cdot M$ (with $N,M$ representing a bipartition), full tomography requires reconstructing $D^{2}-1$ real parameters, while only $D-1$ eigenvalues are needed to characterize the spectrum. Shadow tomography [15], which is restricted to the possible outcomes of the observables to be measured, also scales less favorable than $D-1$. Finally, we note that the AS set is convex, and thus its characterization relies on determining its extremal points.
@@
Several sufficient criteria for AS have been provided over the years. For instance, by bounding matrix norms [16, 17], through linear matrix inequalities [18, 19], by proving the equivalency of absolute separability and absolute PPT for certain subsets [20, 21], using separability witnesses [22] or searching for extremal points [23] among others. Despite these previous achievements, to this date, the set of AS states remains not fully characterized [24].
In this work, we provide a finer geometric characterization of the AS set for arbitrary dimensions. Our results lead to compact analytical conditions of AS which rely on very few or even a single eigenvalue of the state. With similar techniques, we also provide tighter bounds on the geometry of AS set in the multipartite scenario, both for qubits or qudits.
Due to the convexity of the AS set, criteria for AS may be expressed as convex conditions on the state $\rho$. This observation reflects that quantum entanglement, as a resource, cannot be produced (nor amplified) by mere statistical mixing of different separable states. One goal of the present contribution is to join previous criteria with ours into a unifying framework based on convex programs, a class of efficiently solvable optimization algorithms. The separability tests proposed here reconcile the different separability conditions, but also enable to systematically improve over them, for free, without requiring additional experimental effort. In doing so, we bring a new perspective on a long-standing open problem in quantum information theory [8].
@@
The paper is structured as follows: in Section 2, we introduce the formal definitions of AS and its twin problem of absolute positive partial transposition (AP). We review previous criteria, and derive new theorems for AS based on inverting non-completely positive maps. In Section 3, we present our methods to improve present bounds of absolute separability using convex geometrical tools. Importantly, our methods apply to arbitrary dimensions. In Section 4, we address absolute separability in multipartite systems where, again, we derive conditions based on a single eigenvalue alone. Finally, we focus on the symmetric subspace, where analogous conditions for AS and AP can be established. We conclude in Section 5.
@@ 2 Absolute separability
A quantum state is separable if it can be represented as a convex combination of product states. For bipartite states, $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$, (where $\mathcal{B}(\mathcal{H})$ denotes the set of bounded operators acting on $\mathcal{H}$), that implies
$\rho=\sum_{i}p_{i}\ket{e_{i},f_{i}}\bra{e_{i},f_{i}},$
@@
Since unitary transformations preserve the spectrum of a given density matrix $\rho$, it is possible to infer absolute separability criteria based solely on the eigenvalues of the state [20]. For that reason, the set AS is also called separable from spectrum. Clearly, the maximally mixed state (MMS), $\rho\propto\mathds{1}$, is absolutely separable, hence the AS set must contain the MMS and states sufficiently close to it.
The most celebrated criteria for AS is based in the matrix norm of the operator $\rho$, and it is usually referred to as the Gurvits-Barnum ball (GBB) [16]. This criterion can be expressed as a function of the purity of the state, $\mathrm{Tr}(\rho^{2})$, as presented in the following theorem.
@@
From the spectral decomposition of the state $\rho=\sum_{i=0}^{N\cdot M-1}\lambda_{i}\outerproduct{\psi_{i}}{\psi_{i}}$ |, the condition given by Eq. (1) is geometrically interpreted as a ball of radius $(\sqrt{D-1})^{-1}$ in the space of eigenvalues $\bm{\lambda}=\{\lambda_{i}\}$. For normalized states, i.e. $\mathrm{Tr}(\rho)=\sum_{i=0}^{NM-1}\lambda_{i}=1$, this ball is centered on the maximally mixed state.
Absolute separability from linear maps.– Here, we take a different approach, based on linear maps, to tackle the characterization of AS. We recall that positive but not completely positive linear maps, $\Lambda:\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H}^{\prime})$, provide sufficient conditions for entanglement [4, 25, 26] (or, equivalently, necessary conditions for separability) in the form of entanglement witnesses. Interestingly, as reported in [9, 11], such maps are also instrumental to derive sufficient conditions for separability. Before proceeding further, we present the central result on which our work relies on:
@@
In other words, if the set S corresponds to all quantum states acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, and the set $S^{\prime}$ is the subset of separable states, the above theorem provides a sufficient criterion for certifying that a a given state $\sigma\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}$) is separable. As pointed out in [9], the choice of $\Lambda_{\mathbf{p}}$ is determined first by the capability of proving the assumption that $\Lambda_{\bf p}(S)\subset S^{\prime}$, and second by checking that $\Lambda_{\bf p}^{-1}(S^{\prime})\subset S$, that is, by the difficulty of inverting analytically the chosen map.
In this work, we aim at deriving sufficient criteria to detect AS from the previous theorem. To that end, we consider $S$ to be the set of all quantum states, i.e., $\rho\geq 0$ (or, up to normalization, positive-semi definite (PSD) matrices), and $S^{\prime}$ a subset of states which are separable from their spectrum. For this endeavor, it is natural to consider unitarily equivariant linear maps $\Lambda$, i.e., those fulfilling that $\forall\rho,U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\Lambda(U\rho U^{{\dagger}})=U\Lambda(\rho)U^{{\dagger}}$. In this direction, it has been shown [27, 28] that the only unitarily equivariant linear maps are the so-called reduction-like maps [29], which are parameterized as:
@@
with the single parameter $\alpha\in\mathbb{R}$. This family of linear maps is invertible for any nonzero $\alpha$. Its inverse, which is obviously linear, reads:
@@
As demonstrated in [9], the map, $\Lambda_{\alpha}(\rho)$, transforms any state $\rho$ into a separable state for $\alpha\in[-1,2]$. As a result, Theorem 2 shows that the positivity of the inverted map, $\Lambda_{\alpha}^{-1}(\sigma)$ serves as a sufficient condition for separability. By noting that the positivity of $\Lambda_{\alpha}^{-1}(\sigma)$ depends only on the spectrum of state $\sigma$, and that the map is unitarily invariant, it immediately becomes a sufficient condition to certify AS. Moreover, as the conditions for positivity are bounded by the values $\alpha=-1$, and $\alpha=2$, these values lead to sufficient conditions for AS from just a single eigenvalue (either the minimum or maximum).
@@
Let $\rho$ be a normalized bipartite state acting on the space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{N\cdot M+2}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{1}{N\cdot M-1},$ (4)
@@
(See also [9, 30]). For $02\geq\alpha>$ 0, positivity of $\Lambda_{\alpha}^{-1}(\rho)$ demands that $\rho-\mathds{1}/(N\cdot M+\alpha)\geq 0$, that is, $\lambda_{\min}(\rho)-1/(M\cdot N+\alpha)\geq 0$. On the other hand, for $10>\alpha\geq-$ 1, the condition reads $\mathds{1}/(M\cdot N+\alpha)-\rho\geq 0$, or equivalently, $1/(M\cdot N+\alpha)-\lambda_{\max}(\rho)\geq 0$. The extreme values $\alpha=2$ and $\alpha=-1$ yield the announced result. ∎
Finally, let us remark that both, the conditions contained in Theorem 3 and the conditions on Gurvits-Barnum ball, Theorem 1, depend solely on the dimension of the global bipartite system $D=N\cdot M$, independently of the specific local dimensions $N$ and $M$. For a multipartite state of global dimension $D$, they provide sufficient conditions for absolute separability with respect to any bipartition of the total system (see Section 4). As we will show later, the conditions derived from Theorem 3, moreover, provide extremal points in the convex set of the AS set, and permit a geometrical approach to the set of AS states. The characterization of the extreme points of the AS states has been recently achieved but only for specific dimensions; the qubit-qudit and qutrit-qudit case [21, 23].
Recall that all separable states are necessarily positive under partial transposition (PPT), as stated by the famous Peres-Horodecki criterion [3, 31]. It is a natural to ask if there is an analogous set of states that remain positive under partial transposition under the action of any unitary transformations (AP), and if such a set coincides with the set of absolute separable states (AS) [19, 24]. In analogy to the set of AS, the set AP is defined as:
@@
A normalized state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ is absolutely positive under partial transposition (AP) if and only if, for any unitary matrix $U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\rho^{\prime}=U\rho U^{{\dagger}}$ is positive under partial transposition.
In contrast to the AS set, the set of AP has been completely characterized in arbitrary dimensions by a series of linear matrix inequalities (LMI) derived originally in [18]. However, the number of LMI that is necessary to certify AP grows exponentially with $\min\{N,M\}$, which prohibits the characterization for high dimensions. A necessary condition for AP can be derived from the positivity of the first $2\times 2$ block of the first LMI (see Theorem 1 and Section 4 of [18]). Then, considering the eigenvalues of a state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ in non-decreasing order, that is, $\lambda_{0}\leq\lambda_{1}\leq\cdots\leq\lambda_{N\cdot M-1}$, a necessary condition for AP reads as
@@
which involves only the three lowest eigenvalues, $\{\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\}$, and the largest one $\lambda_{N\cdot M-1}$. From Eq. (5) and Theorem 3, and considering the vectors of eigenvalues given by $(1-\frac{D-1}{D+\alpha},\frac{1}{D+\alpha},\cdots,\frac{1}{D+\alpha})$, it is possible to check that only for $\alpha\in[-1,2]$ the necessary condition for AP is fulfilled. Thus, the region for AP coincides with the one for AS.
Remarkably, only for the qubit-qudit case it has been proven that PPT is also sufficient for AS [21]. For arbitrary bipartitions, the equivalence between the AS and AP sets is only conjectured [24]. Although the proposed Theorem 3 does not provide a complete characterization of the AS set, it is valid for any local dimensions and supports the conjecture that AS and AP are identical sets.

### v2

# Sufficient criteria for absolute separability in arbitrary dimensions via linear map inverses

@@ Abstract
Quantum states that remain separable (i.e., not entangled) under any global unitary transformation are known as absolutely separable and form a convex set. Despite extensive efforts, the complete characterization of this set remains largely unknown. In this work, we employ linear maps and their inverses to derive new sufficient analytical conditions for absolute separability in arbitrary dimensions, providing extremal points of this set and improving its characterization. Additionally, we employ convex geometry optimization to refine the characterization of the set when multiple non-comparable criteria for absolute separability are available. We also address the closely related problem of characterizing the absolute PPT (positive partial transposition) set, which consists of quantum states that remain positive under partial transposition across all unitary transformations. Finally, we extend our results to multipartite states. We are proud to dedicate our work to Professor Ryszard Horodecki, whose pioneering contributions to the field of quantum entanglement continue to inspire us all. With deep gratitude and respect.
footnotetext: *These authors contributed equally to this work.footnotetext: $\dagger$} Corresponding author.
@@ 1 Introduction
Despite decades of intense research, necessary and sufficient conditions to ensure separability exist only in the simplest cases [1, 2] or for highly restricted families of states. In fact, entanglement verification and quantification is known to be generically an NP-hard problem [3, 4]. One such necessary and sufficient condition involves positive but not completely positive maps; however, to date, neither a finite set of such maps [5] nor semidefinite programming methods [6, 7] can definitively determine whether a given state is entangled [8, 9]. Given these limitations, rather than seeking iff conditions, one may instead focus on deriving sufficient conditions for separability.
@@
Here, we draw our attention to the use of the inverses of positive but not completely positive maps as a sufficient separability criterion [10, 11, 12]. In particular, we examine absolute separability (AS), which refers to those states that remain separable under all unitary transformations [13]. Characterizing the set of AS is crucial for various reasons. Beyond its significance in fully describing the set of quantum states, AS states play a key role in the resource theory of entanglement, where they represent free states, while entangled states serve as a resource [14, 15]. Since unitary transformations are the primary way of generating entanglement using e.g., quantum circuits or quenches, a state certified as AS cannot become entangled through unitary transformations [16]. Moreover, the characterization of such states must depend only on their spectrum (eigenvalues), and their detection does not require full quantum state tomography [17, 18]. For a quantum state of dimension $D=N\cdot M$ (with $N,M$ representing a bipartition), full tomography requires reconstructing $D^{2}-1$ real parameters, while only D1D-1 eigenvalues are needed to characterize the spectrum. Finally, since the AS set is convex, its characterization relies on identifying its extremal points.
@@
Several sufficient criteria for absolute separability (AS) have been proposed over the years. These include bounding matrix norms [19, 20], using linear matrix inequalities [21, 22], establishing the equivalence of absolute separability and absolute PPT for certain subsets [23, 24], employing separability witnesses [25, 26], computing moments [27], considering positive reduction from the spectrum [28] and identifying extremal points [29], among others. Despite these advancements, the complete characterization of the AS set remains an open problem [30].
Here, we present a more refined characterization of the AS set for arbitrary dimensions. Our findings yield compact analytical conditions for absolute separability that depend on only a few, or even a single eigenvalue of the state. Using similar techniques, we also derive bounds on the AS set in multipartite scenarios, applicable to both qubits and qudits. Due to the convexity of the AS set, criteria for AS can be expressed as convex conditions on the state $\rho$. Here, we show how to merge previous criteria of AS with our own ones into a unifying framework based on convex programming, a class of efficiently solvable optimization algorithms. The separability tests we propose not only reconcile various separability conditions but also enable systematic improvements over them. In doing so, we offer a new perspective on the long-standing open problem of separability from the spectrum in quantum information theory [9].
@@
The paper is organized as follows: In Section 2, we introduce the formal definitions of absolute separability (AS) and its twin problem, absolute positive partial transposition (AP). We review previous criteria and derive new theorems for AS by inverting non-completely positive maps, which significantly improve upon previous results. In Section 3, we present new bounds on absolute separability using convex geometric tools, thereby unifying previously incomparable criteria. Notably, our methods are applicable to arbitrary dimensions. In Section 4, we explore absolute separability in multipartite systems, where we derive analytical conditions based on a single eigenvalue. Finally, we focus on the symmetric subspace, where we establish membership conditions for the AS and AP sets. We conclude in Section 5 and propose several open conjectures stemming from our results.
@@ 2 Absolute separability
Let us recall that a quantum state is separable if it can be represented as a convex combination of product states. For bipartite states, $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$, where $\mathcal{B}(\mathbb{C}^{N})$ denotes the set of bounded operators acting on a Hilbert space $\mathcal{H}=\mathbb{C}^{N}$, that implies
$|,\rho=\sum_{i}p_{i}\ket{e_{i},f_{i}}\bra{e_{i},f_{i}}$,
@@
Since unitary transformations preserve the spectrum of a given density matrix $\rho$, it is possible to infer absolute separability criteria based solely on the eigenvalues of the state [23]. For that reason, the set AS is also called separable from spectrum. Clearly, the maximally mixed state (MMS), $\rho\propto\mathds{1}$, is absolutely separable. Moreover, states sufficiently close to the MMS are also absolutely separable, as discussed in the following.
The most celebrated AS criterion is based on the matrix norm of the operator $\rho$, and is known as the Gurvits-Barnum ball (GBB) [19]. It can be expressed as a function of the purity of the state, $\mathrm{Tr}(\rho^{2})$, as follows:
@@
From the spectral decomposition of the state, $|\rho=\sum_{i=0}^{D-1}\lambda_{i}\outerproduct{\psi_{i}}{\psi_{i}}$, condition Eq. (1) is geometrically interpreted as a ball of radius $(\sqrt{D-1})^{-1}$ in the space of eigenvalues $\bm{\lambda}=\{\lambda_{i}\}$. For normalized states, i.e., $\mathrm{Tr}(\rho)=\sum_{i=0}^{D-1}\lambda_{i}=1$, this ball is centered on the maximally mixed state.
Absolute separability from linear maps.– Here, we take a different approach based on linear maps to tackle the characterization of AS. Recall that positive but not completely positive linear maps, $\Lambda:\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H}^{\prime})$, provide sufficient conditions for entanglement (or, equivalently, necessary conditions for separability) when applied to just one of the subsystems of a bipartite state [2, 31, 32]. Interestingly, as reported in [10, 12], such maps are also instrumental to derive different sufficient conditions for separability. Before proceeding further, we present the central result on which our work relies:
@@
In particular , if the set S corresponds to all quantum states acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, and the set $S^{\prime}$ is the subset of separable states, the above theorem provides a sufficient criterion for certifying that a given state $\sigma\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}$) is separable. As pointed out in [10], the choice of $\Lambda_{\mathbf{p}}$ is determined first by the capability of proving the assumption that $\Lambda_{\bf p}(S)\subset S^{\prime}$, and second by checking that $\Lambda_{\bf p}^{-1}(S^{\prime})\subset S$, that is, by the difficulty of inverting analytically the chosen map.
The previous theorem allows us to derive sufficient criteria for AS. To this end, we define $S$ as the set of all quantum states, i.e., $\rho\geq 0$ (or, up to normalization, positive semi-definite (PSD) matrices), and $S^{\prime}$ as the subset of states that are separable from their spectrum. Within this framework, it is natural to consider unitarily covariant linear maps $\Lambda$, i.e., those fulfilling that for all $\rho$ and $U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\Lambda(U\rho U^{{\dagger}})=V(U)\Lambda(\rho)V(U)^{{\dagger}}$, where $V(U)$ is also unitary [33]. As shown in [34], the only linear maps preserving unitary equivariance on the map and its inverse are the so-called reduction-like maps [35], which are parameterized as:
@@
with the single parameter $\alpha\in\mathbb{R}$. This family of linear maps is invertible for any nonzero $\alpha$. Its inverse, which is linear, reads:
@@
As demonstrated in [10], the map, $\Lambda_{\alpha}(\rho)$, transforms any state $\rho$ into a separable state for $\alpha\in[-1,2]$. Consequently, Theorem 2 establishes that the positivity of the inverted map, $\Lambda_{\alpha}^{-1}(\sigma)$ serves as a sufficient condition for separability. Since the positivity of $\Lambda_{\alpha}^{-1}(\sigma)$ depends only on the spectrum of the state $\sigma$, and the map is unitarily invariant, this condition directly certifies AS. Furthermore, because the conditions for positivity are bounded by $\alpha_{-}=-1$, and $\alpha_{+}=2$, these values provide sufficient criteria for AS based on a single eigenvalue, either the minimum or the maximum, as stated below.
@@
Let $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ be a normalized state in a Hilbert space of global dimension $D=N\cdot M$, with minimum and maximum eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{D+2}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{1}{D-1},$ (4)
@@
(See also [10, 36]). For $2\geq\alpha>0$, positivity of $\Lambda_{\alpha}^{-1}(\rho)$ demands that $\rho-\mathds{1}/(N\cdot M+\alpha)\geq 0$, that is, $\lambda_{\min}(\rho)-1/(M\cdot N+\alpha)\geq 0$. On the other hand, for $0>\alpha\geq-1$, the condition reads $\mathds{1}/(M\cdot N+\alpha)-\rho\geq 0$, or equivalently, $1/(M\cdot N+\alpha)-\lambda_{\max}(\rho)\geq 0$. The extreme values $\alpha=2$ and $\alpha=-1$ yield the announced result. ∎
Finally, we note that both conditions contained in Theorem 3 and the conditions on Gurvits-Barnum ball, Theorem 1, depend solely on the dimension of the global bipartite system $D=N\cdot M$, and are independent of the specific local dimensions $N$ and $M$. For a multipartite state of global dimension $D$, these conditions provide sufficient criteria for absolute separability with respect to any bipartition of the total system (see Section 4). Moreover, as we will show later, the conditions derived from Theorem 3 define extremal points in the convex set of AS states, enabling a geometric approach to its characterization. The classification of the extreme points of AS states is only known for the specific case of qubit-qudit systems [24].
Since all separable states are necessarily positive under partial transposition (PPT), it is natural to ask whether there exists a set of states that remain PPT under any unitary transformation, and whether this set coincides with the AS set [21, 30]. By analogy to the AS set, we define absolute PPT (AP) as:
@@
A normalized state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ is absolutely PPT (AP) if and only if, for any unitary matrix $U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\rho^{\prime}=U\rho U^{{\dagger}}$ is PPT.
Unlike the AS set, the AP set has already been fully characterized for arbitrary dimensions through a series of linear matrix inequalities (LMIs), originally derived in [21]. Despite the number of LMIs required to certify AP grows exponentially with $\min\{N,M\}$, making its characterization infeasible in high dimensions, a necessary AP condition can be obtained from the positivity of the first $2\times 2$ block of the first LMI (see Theorem 1 of [21]). Considering the eigenvalues of a state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ in non-decreasing order, i.e., $\lambda_{0}\leq\lambda_{1}\leq\cdots\leq\lambda_{N\cdot M-1}$, the necessary condition for AP reads:
@@
involving only the three lowest eigenvalues, $\{\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\}$, plus the largest one $\lambda_{N\cdot M-1}$.
Interestingly, the above condition is satisfied only by states whose minimum and maximum eigenvalues fulfill one of the conditions of Theorem 3. That is, those detected with the inverse of the reduction map for parameters $\alpha\in[-1,2]$. Consequently, the region of AP identified by the bounds of $\alpha$ derived from Eq.(5) coincides with the region of AS detected by Eq.(4). Although Theorem 3 does not provide a complete characterization of the AS set, it holds for any local dimensions and in this way, supports the conjecture that AS and AP are identical sets. Only for the qubit-qudit case, $(\mathbb{C}^{2}\otimes\mathbb{C}^{d})$, it has been proven that AP is also a sufficient condition for AS [24]. However, for arbitrary bipartition dimensions, the equivalence between the AS and AP sets remains an open conjecture [30].